OpenStudy (anonymous):

I cannot solve the following first order system of differential equations. I´ve calculated the eigenvalues (λ1 = 1 + i, λ2 = 1 - i) and the eigenvectors E1 = [1, i] , E2 = [i , 1] (column vectors). But, I cannot found a fundamental set of real solutions.

OpenStudy (anonymous):

\[dx/dt = x - y \]\[dy/dt = x + y\] (System of differential equations)

OpenStudy (anonymous):

\[X (t) =\left(\begin{matrix}1 \\ i\end{matrix}\right)e ^{(1+i)t} + \left(\begin{matrix}i \\ 1\end{matrix}\right)e ^{(1-i)t}\] That is the solution, where X(t) is...\[X(t) = \left(\begin{matrix}x(t) \\ y(t)\end{matrix}\right)\] Well, ¿Someone knows how to rewrite it into a set of only real solutions? Please write the steps...

OpenStudy (anonymous):

@mahmit2012 plzzzz help mee http://openstudy.com/study?signup#/updates/4ff05fd6e4b0389b6ee03cd2

OpenStudy (anonymous):